40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog  Not logged in ELOG logo
Message ID: 249     Entry time: Tue Nov 24 12:31:05 2020
Author: Marie K.  
Type: Summary 
Category: BS BRDs 
Subject: LHO BRDs monitoring 

LHO BRDs:

I didn't observe frequency drifts during the month of assembly and monitoring in the Optics Lab. This is not expected from our experiments in the Modal Lab, but it makes the preparation for the sites BRDs easier.

BOUNCE: In anticipation to the frequency drifts, I had tuned the BRDs on the lower side of the target frequency. But the tuning didn't drift so I changed the masses last week for the bounce mode in order to be in the 1% target. The jumps in the curves below are due to retuning (on November 5th for BRD1, on November 20th for the other ones).

Roll: I didn't retune the Roll modes after assembly on October 20th. In the last days, I was experimenting with different ways to excite them (see pictures attached), so this is probably the cause of the slight drifts that we see.

 


 

Attachment 3: Shaker_plate.jpg  3.703 MB  | Hide | Hide all
Shaker_plate.jpg
Attachment 4: Shaker_electronics.jpg  569 kB  | Hide | Hide all
Shaker_electronics.jpg
ELOG V3.1.3-